# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
Understanding the decision-making process of drivers is one of the keys to ensuring road safety. While the driver intent and the resulting ego-motion trajectory is valuable in developing driver-assistance systems, existing methods mostly focus on the motions of other vehicles. In contrast, we focus on inferring the ego-trajectory of a driver’s vehicle using their gaze data. For this purpose, we first collect a new dataset, GEM, which contains high-fidelity ego-motion videos paired with drivers’ eye-tracking data and GPS coordinates. Next, we develop G-MEMP, a novel multimodal ego-trajectory prediction network that combines GPS and video input with gaze data. We also propose a new metric called Path Complexity Index (PCI) to measure the trajectory complexity. We perform extensive evaluations of the proposed method on both GEM and DR(eye)VE, an existing benchmark dataset. The results show that G-MEMP significantly outperforms state-of-the-art methods in both benchmarks. Furthermore, ablation studies demonstrate over 20% improvement in average displacement using gaze data, particularly in challenging driving scenarios with a high PCI. The data, code, and models can be found at https://eth-ait.github.io/g-memp/.
@@ 1 Introduction
Understanding the perception and decision-making process of drivers is crucial for road safety in autonomous and assisted driving. Statistics reveal that 42% of car-bicycle collisions result from driver inattention [1]. In response, a driver-assistance system capable of refining human input based on the predicted intent of the driver would enable vehicles to make informed decisions [38]. While understanding drivers’ intent is critical for inferring their future trajectories, traditional methods mostly focus on predicting the paths of other vehicles [36, 13, 25, 46, 26]. In contrast, we seek to predict drivers’ ego-trajectory and to infer their intent with the help of gaze input.
Figure 1: G-MEMP framework. Using the scene videos, gaze, and past GPS coordinates, we predict the future trajectory (in comparison to the target trajectory). The training is regularized through predicted scene and gaze features.
This setting presents a dynamic interplay of imminent and long-term intent, influenced by the surrounding traffic. Our central idea builds on the insight that the gaze of a person is tightly intertwined with their intent in varying horizons [53, 15, 3]. Especially in driving, the direction and focus of a driver’s gaze can reveal important information about their decisions on the road. For instance, a driver whose gaze fixates on the brake lights of the vehicle ahead is likely to decelerate promptly, reflecting an imminent intent to slow down, as seen in Figure 2. While regular driving scenarios range from steady cruising to navigating intersections, the greatest value of being able to predict ego-trajectories emerge in ambiguous, complex road paths, e.g., cases where the road is not trivially straight or when predicting where the driver would turn at an intersection. Our focus is thus twofold: firstly, to develop an efficient method for detecting these rare yet valuable scenarios in existing data, and secondly, to establish a framework capable of understanding the driver’s intent in such complex situations for ego-trajectory prediction. To this end, we present a new dataset, a novel metric, and an end-to-end framework.
@@
We first introduce a Gaze-assisted Ego Motion (GEM) dataset that captures diverse urban driving scenarios across 10 subjects, integrating GPS, eye gaze, and high-resolution scene data from three cameras—two frontal dash cameras and one positioned on the driver’s glasses (see Figure 1). To bolster accuracy, GPS data has been manually corrected, ensuring precise capture of vehicular speed and acceleration dynamics. To better quantify the task complexity, we propose a new metric named Path Complexity Index (PCI), which measures the divergence of a driving trajectory from an anticipated path. It allows us to link trajectory complexity with model performance. Finally, we build a multi-modal framework, G-MEMP (Gaze-Enhanced Multimodal Ego-Motion Prediction), for egocentric trajectory estimation that integrates scene data, gaze patterns, and past trajectory (see Figure 1). We train G-MEMP with auxiliary tasks of predicting future gaze and scene features for regularization, using a novel future-discounted loss formulation. Both qualitative and quantitative experimental results show that G-MEMP outperforms state-of-the-art (SOTA) methods on the GEM dataset. To test the generalization ability, we also evaluate G-MEMP on the DR(eye)VE dataset [31], and our method surpasses the SOTA approaches. Ablation studies show that incorporating the gaze information improves the prediction quality, and our proposed model particularly excels in high-PCI situations.
@@
  • GEM, a dataset with high-fidelity ego-motion videos coupled with accurate GPS coordinates and gaze data.

### v2

@@ Abstract
Understanding drivers’ decision-making is crucial for road safety. Although predicting the ego-vehicle’s path is valuable for driver-assistance systems, existing methods mainly focus on external factors like other vehicles’ motions, often neglecting the driver’s attention and intent. To address this gap, we infer the ego-trajectory by integrating the driver’s gaze and the surrounding scene. We introduce RouteFormer, a novel multimodal ego-trajectory prediction network combining GPS data, environmental context, and driver field-of-view—comprising first-person video and gaze fixations. We also present the Path Complexity Index (PCI), a new metric for trajectory complexity that enables a more nuanced evaluation of challenging scenarios. To tackle data scarcity and enhance diversity, we introduce GEM, a comprehensive dataset of urban driving scenarios enriched with synchronized driver field-of-view and gaze data. Extensive evaluations on GEM and DR(eye)VE demonstrate that RouteFormer significantly outperforms state-of-the-art methods, achieving notable improvements in prediction accuracy across diverse conditions. Ablation studies reveal that incorporating driver field-of-view data yields significantly better average displacement error, especially in challenging scenarios with high PCI scores, underscoring the importance of modeling driver attention. All data and code is available at meakbiyik.github.io/routeformer.
@@ 1 Introduction
Understanding the perception and decision-making process of drivers is crucial for road safety in autonomous and assisted driving. Statistics reveal that 42% of car-bicycle collisions result from driver inattention (Allenbach et al., 2021). In response, a driver-assistance system capable of integrating the driver’s perspective into its decision-making would enable vehicles to make informed decisions (Schwarting et al., 2018). Although predicting the driver’s path is critical for such systems, traditional methods focus mainly on inferring the paths of other vehicles (Salzmann et al., 2020; Gu et al., 2021; Ngiam et al., 2021; Varadarajan et al., 2022; Nayakanti et al., 2023). In contrast, we seek to predict the drivers’ ego-trajectory by observing both the surrounding scene, and the perception of the ego-vehicle’s drivers via their frontal view and gaze fixations.
The central idea of this work builds on the insight that the perception of a person is tightly intertwined with their imminent and long-term goals, influenced by the surrounding traffic (Triesch et al., 2003; Hayhoe et al., 2003; Argyle et al., 1994). In driving, the attention of a driver can reveal otherwise inaccessible information about their decisions on the road. Complementing scene information with these cues enables the vehicles to resolve ambiguous situations. For instance, the head movements of the driver might be highly informative in determining the turn direction in a junction. The value of predicting ego-trajectories emerge in ambiguous, complex road paths, e.g., cases where the road is not trivially straight, or when predicting where the driver would turn at an intersection.
Our focus is thus twofold: firstly, to develop a framework capable of predicting the drivers’ future ego-trajectories in various driving scenarios, and secondly, to create a method for identifying complex yet rare trajectories and quantifying their complexity. Despite the presence of other driving datasets with gaze data (Gopinath et al., 2021; Xia et al., 2017; Fang et al., 2023), to the best of our knowledge, only one public benchmark DR(eye)VE contains both ego-drivers’ locations and their field-of-view (Palazzi et al., 2018). The scarcity of available data poses another challenge for our task. To this end, we present an end-to-end framework that predicts future ego-trajectories, a novel metric to measure the complexity of road scenarios, and a new dataset of real-world driving recordings that include drivers’ field-of-view changes.
Figure 1: RouteFormer framework. Using the past GPS and the scene together with driver field-of-view, we predict the future ego-trajectory and visual features in driving with a novel loss scheme.
@@
First, we build a multi-modal framework, RouteFormer, for egocentric trajectory estimation that integrates scene data, driver FOV, and past trajectory, as shown in Figure 1. RouteFormer is crafted based on insights from time-series forecasting literature, offering the flexibility to incorporate additional data from other modalities, such as navigation paths and LIDAR. We train RouteFormer with auxiliary tasks of predicting future gaze and scene features for regularization, using a novel future-discounted loss formulation. To better quantify the complexity of various driving scenarios, we propose a new metric named Path Complexity Index (PCI), which measures the divergence of a driving trajectory from an extrapolation of the current path. It also indicates the difficulty of the corresponding prediction tasks and allows us to analyze model performance under different scenarios. Finally, we introduce a Gaze-assisted Ego Motion (GEM) dataset that captures diverse urban driving scenarios across 10 subjects, integrating GPS, high-resolution scene data from two front-view cameras, and driver field-of-view (FOV)—comprising first-person video and gaze fixations.
Experimental results show that RouteFormer outperforms state-of-the-art (SOTA) methods on the GEM dataset. To test its generalization capabilities, we also evaluate RouteFormer on the DR(eye)VE dataset (Palazzi et al., 2018), and our method surpasses the SOTA approaches. Ablation studies show that incorporating drivers’ field-of-view information improves the prediction quality, particularly in high-PCI situations.
@@
  • RouteFormer, an end-to-end multimodal ego-motion prediction network that effectively utilizes FOV data with a novel loss design.